### v1

@@ 1 Introduction
Figure 1: Schematic configurations. We take the motivating laboratory experiment on the left, create its digital twin, and subsequently simplify the latter successively. Identifiers for different configurations are included above each sketch. The location of different temperatures is defined on the left, whereas panels (b – f) include the corresponding control parameters only; other values manifest dynamically.
@@
Of course, these limits can only be approximated by real materials and material selection is often further constrained by requirements of the measurement techniques such as optical transparency. Figure 1 (a) shows such a compromise when analysing the large-scale heat transfer patterns (Moller et al., 2020; Moller, 2022). On the one hand, the glass plate on top of the fluid layer of interest allows optical access, but it leads to $\kappa_{\textrm{st}}/\kappa_{\textrm{fl}}=2.5$ and needs to be cooled using a pressure-driven flow. On the other hand, the aluminium bottom plate offers $\kappa_{\textrm{sb}}/\kappa_{\textrm{fl}}=435$ and can be heated quite uniformly via meandering channels. Despite a small temperature difference of approximately $2K$ between both plates (Moller et al., 2022; Moller, 2022), a convection flow emerges in the water which is visualised by suspended (temperature-sensitive) tracer particles in a horizontal slab at mid-height. This asymmetry of thermal boundary conditions has so far not been accounted for in simulations, resulting in unresolved discrepancies when comparing experimental and numerical results. Most strikingly, experiments suggested an increased size of the turbulent superstructures and decreased induced heat transfer across the fluid layer compared to simulations (Pandey et al., 2018; Moller et al., 2021, 2022; Vieweg et al., 2021, 2021a; Schneide et al., 2022; Käufer et al., 2023). Unfortunately, precise attributions of those disagreements have not yet been possible due to a lack of simulations that resemble the experiment – especially with respect to its thermal boundary conditions, strength of thermal driving, and horizontal extent of the domain; all of which affect those disagreements – sufficiently.
@@ 2.2 Numerical domain and its boundary conditions
In the iterative process of simplifying the configuration, see again figure 1, we might eventually substitute this Newton cooling boundary condition from equation (2.9b) by another Dirichlet condition
@@ 3.2 Simplification of the numerical domain
For these reasons, we study the effect of varying thermal boundary conditions by considering different configurations of the numerical domain – in terms of both the thermophysical and geometric properties – as presented in figure 1. Commencing with the digital twin, we successively simplify the configuration. First, we replace the Newton cooling boundary condition (2.9b) by another (stronger) Dirichlet condition (2.11) (see panel (c)). Second, we replace the glass top plate by an aluminium one such that $\kappa_{\textrm{st}}/\kappa_{\textrm{fl}}=\kappa_{\textrm{sb}}/\kappa_{\textrm{fl}}$ (see panel (d)). Third, the thickness of the top plate is adjusted to that of the bottom plate such that $\Gamma_{\textrm{st}}=\Gamma_{\textrm{sb}}$ (see panel (e)). Fourth and finally, we omit the solid top and bottom aluminium plates entirely and apply Dirichlet conditions directly to the fluid layer (see panel (f)) – this situation corresponds to $\left\{\Gamma_{\textrm{st}},\Gamma_{\textrm{sb}}\right\}\rightarrow 0$ and converged to or represents the traditional Rayleigh-Bénard convection configuration. Note that the different successive modifications build up on each other.
@@ 3.3 Comparison of different configurations of the numerical domain
Table 2: Global characteristic measures of the simulations from table 1. This table contains the maximum instantaneous temperature difference at the upper solid-fluid interface $\max\left(\Delta_{\textrm{hor}}T_{\textrm{t}}\right)$, the instantaneous standard deviation of the temperature field at this interface $\textrm{std}\left(T_{\textrm{t}}\right)$, the true global Nusselt number $\mathrm{Nu}$ (which includes the diffusive heat transport), the experimentally accessible Nusselt number $\mathrm{Nu}_{\textrm{exp}}$, the Reynolds number $\mathrm{Re}$, as well as the integral length scale of the temperature field $\Lambda_{T}$. All values are provided as temporal means together with the corresponding standard deviation. Motivating laboratory experiments are contrasted via rows with grey text colour and are based on a restricted field.
@@ 4.4 Re-assessment of the original laboratory measurement data
We therefore carefully scrutinise or re-assess the processing of the original stereo-PIV measurement data starting from the raw uncalibrated camera images. This reveals that the previously used pinhole camera calibration model with a subsequent self-calibration (Wieneke, 2005) did not account for the optical refraction effect between the cameras and the fluid layer, resulting in wrong vertical distances of the calibration planes and thus a systematic underestimation of only the vertical or out-of-plane velocity component. We thus re-process the PIV data using the better suited polynomial calibration model and find a relative increase of $u_{z}$ by approximately $24\%$ for all different Rayleigh numbers. As this number propagates directly to $\mathrm{Nu}_{\textrm{exp}}$, they agree now almost perfectly with the expectations based on the digital twin. This re-assessment reveals new (corrected) global measures of heat and momentum transport in the laboratory experiment of $\mathrm{Nu}_{\textrm{exp}}=\left\{5.09,7.04,7.25\right\}$ and $\mathrm{Re}=\left\{13.56,20.99,28.55\right\}$, respectively. Figure 6 (g – l) includes also the corrected statistical data.
@@ 5 Discussion and perspective
We find for this twin that $\max\left(\Delta_{\textrm{hor}}T_{\textrm{t}}\right)>0.5$ and thus exceeds half of the total temperature drop $\Delta T$ across the fluid layer for all $\mathrm{Ra}$. This shows that the upper solid-fluid interface is extremely prone to thermal inhomogeneities and proves that the past assumption ‘the isothermal boundary conditions can be considered as fulfilled in good approximation’ (Moller et al., 2020) has to be reconsidered. Only the bottom aluminium plate renders the corresponding interface iso-thermal. A comparison of the different convection and heat transfer coefficients $\left\{h_{\mathrm{Bi}},h_{\textrm{st}},h_{\textrm{fl}},h_{\textrm{sb}}\right\}$ across the different layers (see again equations (2.15d) and (2.15e)) shows that thermal conduction across the top plate represents the essential bottleneck for heat transfer in the domain and thus alters the boundary conditions significantly. For instance, $h_{\textrm{st}}/h_{\textrm{fl}}\approx 1.6$ in contrast to $h_{\mathrm{Bi}}/h_{\textrm{fl}}\approx 2.9$ and $h_{\textrm{sb}}/h_{\textrm{fl}}\approx 126$ for $\mathrm{Ra}=2\times 10^{5}$. As is underlined by simulation 2CHTa (corresponding to $h_{\mathrm{Bi}}\rightarrow\infty$), the effect of increasing $\mathrm{Bi}$ is very limited.
@@
Eventually, the comparison of the vertical velocities $u_{z}$ from both the experiment and its digital twin suggested a systematic underestimation in the case of the former. In fact, this out-of-plane velocity component is most susceptible to systematic measurement errors during stereoscopic PIV (Westerweel, 1997; Prasad, 2000; Cierpka et al., 2012; Raffel et al., 2018) with an exact calibration being key (Prasad, 2000; Wieneke, 2005). A subsequent re-assessment of the original stereo-PIV measurement data starting from the raw camera images revealed indeed a camera calibration error in (Moller et al., 2021, 2022; Moller, 2022), underestimating the distances between the calibration planes. Correcting this mistake results in a $24\%$ relative increase of $u_{z}$ – the sign and distribution is not affected. This allows to finally to collapse the data from both the laboratory experiment as well as the numerical simulations across the entire range of considered $\mathrm{Ra}$, particularly with respect to $\mathrm{Nu}$. This resolves the remaining motivating discrepancies.
This study highlights that digital twins represent, together with the resemblance of laboratory measurement procedures, a highly useful tool for resolving discrepancies between experimental and numerical observations and thus drive the progress in thermofluid science with its numerous applications. From an experimental perspective, our study suggests moving towards volumetric Lagrangian particle tracking techniques to prevent incorrect reconstructions of individual velocity components (Käufer and Cierpka, 2024). Especially in combination with physics-informed machine learning, this allows revealing even more information of the flow (Toscano et al., 2024). From a physical perspective, our study underlines the crucial role of realistic thermal boundary conditions with respect to the formation of long-living large-scale flow structures as well as their characteristic size and lifetime. It is clear that this point becomes even more important when the geometry of the heat transfer system goes beyond a simple cuboid configuration as discussed here. Understanding the effect of symmetric non-ideal thermal boundary conditions is essential for a successful interpretation of more complex configurations and will be addressed in a future study.

### v2

@@ 1 Introduction
Figure 1: Schematic configurations. We take the motivating laboratory experiment (EXP) on the left, create its digital twin which involves a Newton cooling (NC) condition, and subsequently simplify the latter successively. Identifiers for different configurations are included above each sketch. The location of different temperatures is defined on the left, whereas panels (b – f) include the corresponding control parameters only; other values manifest dynamically. CHT stands for regular conjugate heat transfer and DIR for pure or classical Dirichlet boundary conditions.
@@
Of course, these limits can only be approximated by real materials and material selection is often further constrained by requirements of the measurement techniques such as optical transparency. Figure 1 (a) shows such a compromise when analysing the large-scale heat transfer patterns (Moller et al., 2020; Moller, 2022). On the one hand, the glass plate on top of the fluid layer of interest allows optical access, but it leads to $\kappa_{\textrm{st}}/\kappa_{\textrm{fl}}=2.5$ and needs to be cooled using a pressure-driven flow. On the other hand, the aluminium bottom plate offers $\kappa_{\textrm{sb}}/\kappa_{\textrm{fl}}=435$ and can be heated quite uniformly via meandering channels. Despite a small temperature difference of approximately $2K$ between both plates (Moller et al., 2022; Moller, 2022), a convection flow emerges in the water which is visualised by suspended (temperature-sensitive) tracer particles in a horizontal slab at mid-height. A detailed description of the laboratory experiment can be found in Moller (2022). This asymmetry of thermal boundary conditions has so far not been accounted for in simulations, resulting in unresolved discrepancies when comparing experimental and numerical results. Most strikingly, experiments suggested an increased size of the turbulent superstructures and decreased induced heat transfer across the fluid layer compared to simulations (Pandey et al., 2018; Moller et al., 2021, 2022; Vieweg et al., 2021, 2021a; Schneide et al., 2022; Käufer et al., 2023). Unfortunately, precise attributions of those disagreements have not yet been possible due to a lack of simulations that resemble the experiment – especially with respect to its thermal boundary conditions, strength of thermal driving, and horizontal extent of the domain; all of which affect those disagreements – sufficiently.
@@ 2.2 Numerical domain and its boundary conditions
In the iterative process of simplifying the configuration, see again figure 1, we might eventually substitute this Newton cooling (NC) boundary condition from equation (2.9b) by another Dirichlet condition
@@ 3.1 The digital twin
Note that while thermal conduction through or in the lateral walls can become significant especially for slender convection cells (Stevens et al., 2014), we assume them to be adiabatic in our motivating experimental set-up (and thus also the digital twin, see eq. (2.7)) due to several reasons. First, the experiment offers a large aspect ratio $\Gamma\gg 1$ and so the total area of the side walls is by an order of magnitude smaller than that of the top and bottom plates. Second, the mean temperature in the turbulent flow is held at room temperature during experiments and the side walls are additionally insulated. And third, both the fluid’s heat transfer coefficient and horizontal cross section are significantly larger than those of the side walls, rendering vertical heat transfer within the latter negligible.
@@ 3.2 Simplification of the numerical domain
For these reasons, we study the effect of varying thermal boundary conditions by considering different configurations of the numerical domain – in terms of both the thermophysical and geometric properties – as presented in figure 1. Commencing with the digital twin, we successively simplify the configuration. First, we replace the Newton cooling (NC) boundary condition (2.9b) by another (stronger) Dirichlet condition (2.11) (see panel (c)) similar to typical conjugate heat transfer (CHT) problems (Perelman, 1961; Foroozani et al., 2021). Second, we replace the glass top plate by an aluminium one such that $\kappa_{\textrm{st}}/\kappa_{\textrm{fl}}=\kappa_{\textrm{sb}}/\kappa_{\textrm{fl}}$ (see panel (d)). Third, the thickness of the top plate is adjusted to that of the bottom plate such that $\Gamma_{\textrm{st}}=\Gamma_{\textrm{sb}}$ (see panel (e)). Fourth and finally, we omit the solid top and bottom aluminium plates entirely and apply Dirichlet conditions (DIR) directly to the fluid layer (see panel (f)) – this situation corresponds to $\left\{\Gamma_{\textrm{st}},\Gamma_{\textrm{sb}}\right\}\rightarrow 0$ and converged to or represents the traditional Rayleigh-Bénard convection configuration. Note that the different successive modifications build up on each other.
@@ 3.3 Comparison of different configurations of the numerical domain
Table 2: Global characteristic measures of the simulations from table 1. This table contains the maximum instantaneous temperature difference at the upper solid-fluid interface $\max\left(\Delta_{\textrm{hor}}T_{\textrm{t}}\right)$, the instantaneous standard deviation of the temperature field at this interface $\textrm{std}\left(T_{\textrm{t}}\right)$, the true global Nusselt number $\mathrm{Nu}$ (which includes the diffusive heat transport), the experimentally accessible Nusselt number $\mathrm{Nu}_{\textrm{exp}}$, the Reynolds number $\mathrm{Re}$, as well as the integral length scale of the temperature field $\Lambda_{T}$. All values are provided as temporal means together with the corresponding standard deviation. Motivating laboratory experiments are contrasted via rows with grey text colour and are based on a restricted field. Revised values of their $\mathrm{Nu}_{\textrm{exp}}$ and $\mathrm{Re}$ are reported in section 4.4.
@@ 4.4 Re-assessment of the original laboratory measurement data
We therefore carefully scrutinise or re-assess the processing of the original stereo-PIV measurement data starting from the raw uncalibrated camera images. This reveals that the previously used pinhole camera calibration model with a subsequent self-calibration (Wieneke, 2005) did not account for the optical refraction effect between the cameras and the fluid layer, resulting in wrong vertical distances of the calibration planes and thus a systematic underestimation of only the vertical or out-of-plane velocity component. This issue can be eliminated by using a polynomial calibration which creates the image-to-world mapping by fitting polynomials through the calibration markers. This approach does not require any modeling of the optical path and is well-suited for complex set-ups. As a downside, it allows for limited extrapolations beyond the calibrated region only which is not of relevance for us anyway. We thus re-process the PIV data using a polynomial calibration and find a relative increase of $u_{z}$ by approximately $24\%$ for all different Rayleigh numbers. As this number propagates directly to $\mathrm{Nu}_{\textrm{exp}}$, they agree now almost perfectly with the expectations based on the digital twin. This re-assessment reveals new (corrected) global measures of heat and momentum transport in the laboratory experiment of $\mathrm{Nu}_{\textrm{exp}}=\left\{5.09,7.04,7.25\right\}$ and $\mathrm{Re}=\left\{13.56,20.99,28.55\right\}$, respectively. Figure 6 (g – l) includes also the corrected statistical data.
@@ 5 Discussion and perspective
We find for this twin that $\max\left(\Delta_{\textrm{hor}}T_{\textrm{t}}\right)>0.5$ and thus exceeds half of the total temperature drop $\Delta T$ across the fluid layer for all $\mathrm{Ra}$. This shows that the upper solid-fluid interface is extremely prone to thermal inhomogeneities and proves that the past assumption ‘the isothermal boundary conditions can be considered as fulfilled in good approximation’ (Moller et al., 2020) has to be reconsidered for the experimentally covered range of $\mathrm{Ra}$. Only the bottom aluminium plate renders the corresponding interface iso-thermal. A comparison of the different convection and heat transfer coefficients $\left\{h_{\mathrm{Bi}},h_{\textrm{st}},h_{\textrm{fl}},h_{\textrm{sb}}\right\}$ across the different layers (see again equations (2.15d) and (2.15e)) shows that thermal conduction across the top plate represents the essential bottleneck for heat transfer in the domain and thus alters the boundary conditions significantly. For instance, $h_{\textrm{st}}/h_{\textrm{fl}}\approx 1.6$ in contrast to $h_{\mathrm{Bi}}/h_{\textrm{fl}}\approx 2.9$ and $h_{\textrm{sb}}/h_{\textrm{fl}}\approx 126$ for $\mathrm{Ra}=2\times 10^{5}$. As is underlined by simulation 2CHTa (corresponding to $h_{\mathrm{Bi}}\rightarrow\infty$), the effect of increasing $\mathrm{Bi}$ is very limited.
@@
Eventually, the comparison of the vertical velocities $u_{z}$ from both the experiment and its digital twin suggested a systematic underestimation in the case of the former. In fact, this out-of-plane velocity component is most susceptible to systematic measurement errors during stereoscopic PIV (Westerweel, 1997; Prasad, 2000; Cierpka et al., 2012; Raffel et al., 2018) with an exact calibration being key (Prasad, 2000; Wieneke, 2005). A subsequent re-assessment of the original stereo-PIV measurement data starting from the raw camera images revealed indeed a camera calibration error in (Moller et al., 2021, 2022; Moller, 2022) and thus also the data used in (Käufer et al., 2023; Teutsch et al., 2023), underestimating the distances between the calibration planes. Correcting this mistake results in a $24\%$ relative increase of $u_{z}$ – the latter’s sign and distribution, as well as the size and temporal evolution of flow structures, are not affected. This allows to finally to collapse the data from both the laboratory experiment as well as the numerical simulations across the entire range of considered $\mathrm{Ra}$, particularly with respect to $\mathrm{Nu}$. This resolves the remaining motivating discrepancies.
This study highlights that digital twins represent, together with the resemblance of laboratory measurement procedures, a highly useful tool for resolving discrepancies between experimental and numerical observations and thus drive the progress in thermofluid science with its numerous applications. From an experimental perspective, our study suggests moving towards volumetric Lagrangian particle tracking techniques to prevent incorrect reconstructions of individual velocity components (Käufer and Cierpka, 2024). Especially in combination with physics-informed machine learning, this allows revealing even more information of the flow (Toscano et al., 2024). From a physical perspective, our study underlines the crucial role of realistic thermal boundary conditions with respect to the formation of long-living large-scale flow structures as well as their characteristic size and lifetime. It is clear that this point becomes even more important when the geometry of the heat transfer system goes beyond a simple cuboid configuration as discussed here. Understanding the effect of symmetric non-ideal thermal boundary conditions is essential for a successful interpretation of more complex configurations and will be addressed in a future study. Another point that has to be left open for future work is the impact of these non-ideal boundary effects at higher Rayleigh numbers.